OpenStudy (anonymous):

Use the distributive property to remove the parentheses. (y+9)11

OpenStudy (whpalmer4):

The distributive property says \[a(b+c) = a*b + a*c\]and because of the commutative property, that means \[(b+c)a = a*b + a*c\]as well

OpenStudy (anonymous):

okay so.........

OpenStudy (anonymous):

Distribute the term 11 by multiplying to both terms 11(y+9) 11y+99

OpenStudy (anonymous):

@theonlycool13

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=distribute+11%28y%2B9%29 Here's the proof if you don't believe me, also use this website to double-check :)

OpenStudy (anonymous):

Any q's?

OpenStudy (anonymous):

it was wrong thanks any way

OpenStudy (anonymous):

My answer?? What was the correct one? Last time I checked the distributive property, is pretty straightforward......

OpenStudy (anonymous):

okay sorry but it was wrong

OpenStudy (anonymous):

No need to be sorry! I'd like the name of that teacher the distributive prop, in this case is simple multiplication! Like @whpalmer4 posted!!! thats a mathematical rule :P You should definitely protest answer

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what was the right answer....

OpenStudy (anonymous):

cant it is a virtual school it online

OpenStudy (anonymous):

11y+99

OpenStudy (anonymous):

That's the one i posted above.... here all copy/ paste no changes..... Distribute the term 11 by multiplying to both terms 11(y+9) 11y+99

OpenStudy (anonymous):

The link i provided arrived to the same conclusion i did, if u click link :)

OpenStudy (whpalmer4):

yes, there's no question that \[(y+9)11 = 11y + 99 = 99 + 11y\]

OpenStudy (whpalmer4):

Why are you attending a virtual school?

OpenStudy (anonymous):

so sorry i thought it was 11(y+9)

OpenStudy (anonymous):

how was my answer wrong.... grr frustrating

OpenStudy (anonymous):

ohh sorry should have made more clear... I thought you would follow my steps

OpenStudy (anonymous):

no ur anser was right

OpenStudy (anonymous):

oh i am so stupid

OpenStudy (anonymous):

so sorry

OpenStudy (anonymous):

no, no not stupid! I should have confirmed you understood, half my bad :)

OpenStudy (anonymous):

this stuff is hard for me i need lots of help can you?

OpenStudy (anonymous):

no yo cool

OpenStudy (anonymous):

what do u need help w/?

OpenStudy (anonymous):

the same kind of problem 5 in fact

OpenStudy (anonymous):

go ahead and post, or u can close this q and post another and post my @Nanalew, so i can just link to it either way :)

OpenStudy (anonymous):

K will :-)

OpenStudy (anonymous):

10(v-3) @Nanalew

OpenStudy (anonymous):

10v-30 :) You can double-check w/ the website i used earlier :)

OpenStudy (anonymous):

Understand :) All u r doing is multiplying 10 to both terms :) Or in other terms distributing

OpenStudy (anonymous):

i hav to eat

OpenStudy (anonymous):

okay back